### v1

@@ 3.6 Supernova contamination
Our observations are sensitive to the presence of stellar transients, especially supernovae (SNe), which may be a source of contamination in our counting of AGN. In Paper I we identified three variables sources that are supernovae, two of which are apparently hostless with indeterminable redshifts. Supernovae are expected to follow either the stellar mass distribution (thermonuclear) or star formation (SFR) distribution (core collapse). Statistically in a galaxy these are centrally concentrated, so our approach of focussing on deep-stack centroiding in galactic nuclei, while helping to avoid off-nuclear SNe, would still be subject to some level of stellar transient contaminants.
@@ 4 Conclusions
Historically, high-redshift luminous quasars have been known for some time (Fan et al., 2006; Mortlock et al., 2011; Bañados et al., 2018), though they are typically much brighter—by four magnitudes or more—than the sources we detect. More recently, there have been reports of AGN at even higher redshifts with luminosities closer to $L^{\star}$ (e.g., Maiolino et al., 2023; Harikane et al., 2023; Larson et al., 2023). However, our survey, which was restricted to the small volume of the HUDF, has the ability to place stronger constraints than any previous work on the number densities of SMBHs at these redshifts. We have shown how the 25 variable sources we detect between $z=6$ and 9 imply high co-moving number densities of $n_{\rm SMBH}\gtrsim 10^{-2}\>{\rm cMpc}^{-3}$. In addition, our results at low redshift make a new measurement of $n_{\rm SMBH}\sim 6\times 10^{-2}\>{\rm cMpc}^{-3}$ in this regime that is about a factor of 10 higher than previous estimates.
Our measurements constrain SMBH seeding mechanisms and their implementation in cosmological simulations. For example, our estimate of $n_{\rm SMBH}$ at $z=7-9$ is about a factor of 50 higher than expected in the fiducial halo mass threshold model used in the Illustris family of simulations (Vogelsberger et al., 2014). Our estimates, including at $z\sim 0$, are $\sim 10^{2}-10^{4}$ times larger than produced in simulations of SMBH seeding via Direct Collapse (Chon et al., 2016; Wise et al., 2019; Latif et al., 2022).
@@ Acknowledgements
We thank Yuichi Harikane and Anna Wright for helpful discussions. We thank Peter Williams for providing the 3-color image of the HUDF used in Fig. 4 and scripts for overplotting data. V.C. acknowledges support from the Chalmers Initiative on Cosmic Origins (CICO) as a visiting Ph.D. student at Chalmers Univ. and co-funding of his Ph.D. position at Univ. of Trieste. J.C.T. acknowledges support from ERC Advanced grant 788829 (MSTAR). A.R.Y. is supported by the Swedish National Space Agency (SNSA). M.J.H. is supported by the Swedish Research Council (Vetenskapsrådet) and is Fellow of the Knut & Alice Wallenberg Foundation. R.S.E. acknowledges generous financial support from the Peter and Patricia Gruber Foundation. M.J.H. and J.C.T. thank the staff in The Doors public house for service that germinated the ideas that led to this project.
@@
The inclusion of a Data Availability Statement is a requirement. Data Availability Statements provide a standardised format for readers to understand the availability of data underlying the research results described in the article. The statement may refer to original data generated in the course of the study or to third-party data analysed in the article. The statement should describe and provide means of access, where possible, by linking to the data or providing the required accession numbers for the relevant databases or DOIs. All the HST data used in this paper can be found in MAST: http://dx.doi.org/10.17909/7s5v-gz68 (catalog 10.17909/7s5v-gz68).
@@ Appendix A Variability in F105W and F160W
Figures 7 and 8 report the photometric variability as a function of the mean mag in the F105W and F160W filters, as detailed in Fig. 1.

### v2

@@ 3.6 Supernova contamination
Our observations are sensitive to the presence of stellar transients, especially supernovae (SNe), which may be a source of contamination in our counting of AGN. In Paper I we identified three variables sources that are supernovae, two of which are apparently hostless with indeterminable redshifts. Supernovae are expected to follow either the stellar mass distribution (thermonuclear) or star formation (SFR) distribution (core collapse). Statistically in a galaxy these are centrally concentrated, so our approach of focussing on deep-stack centroiding in galactic nuclei, while helping to avoid off-nuclear SNe, would still be subject to some level of stellar transient contaminants. We note that we do not consider Tidal Disruption Events (TDEs), i.e, stellar disruption by a supermassive black hole, to be contaminants in our census.
@@ 4 Conclusions
Historically, high-redshift luminous quasars have been known for some time (Fan et al., 2006; Mortlock et al., 2011; Bañados et al., 2018), though they are typically much brighter—by four magnitudes or more—than the sources we detect. More recently, there have been reports of AGN at even higher redshifts with luminosities closer to $L^{\star}$ (e.g., Maiolino et al., 2023; Harikane et al., 2023; Larson et al., 2023). However, our survey, which was restricted to the small volume of the HUDF, has the ability to place stronger constraints than any previous work on the number densities of SMBHs at these redshifts. We have shown how the 13 variable sources we detect between $z=6$ and 8 imply high co-moving number densities of $n_{\rm SMBH}\gtrsim 2\times 10^{-3}\>{\rm cMpc}^{-3}$. In addition, our results at low redshift make a new measurement of $n_{\rm SMBH}\sim 2\times 10^{-2}\>{\rm cMpc}^{-3}$ in this regime that is about a factor of 5 higher than previous estimates.
Our measurements constrain SMBH seeding mechanisms and their implementation in cosmological simulations. For example, our estimate of $n_{\rm SMBH}$ at z=78z=7-8 is about a factor of 10 higher than expected in the fiducial halo mass threshold model used in the Illustris family of simulations (Vogelsberger et al., 2014). Our estimates, including at $z\sim 0$, are $\sim 10^{2}-10^{4}$ times larger than produced in simulations of SMBH seeding via Direct Collapse (Chon et al., 2016; Wise et al., 2019; Latif et al., 2022).
@@ Acknowledgements
We thank Yuichi Harikane and Anna Wright for helpful discussions. We thank Peter Williams for providing the 3-color image of the HUDF used in Fig. 4 and scripts for overplotting data. V.C. thanks the BlackHoleWeather project and PI Prof. Gaspari for salary support. V.C. also acknowledges support from the Chalmers Initiative on Cosmic Origins (CICO) as a visiting Ph.D. student at Chalmers Univ. and co-funding of his former Ph.D. position at Univ. of Trieste. J.C.T. acknowledges support from ERC Advanced grant 788829 (MSTAR). A.R.Y. is supported by the Swedish National Space Agency (SNSA). M.J.H. is supported by the Swedish Research Council (Vetenskapsrådet) and is Fellow of the Knut & Alice Wallenberg Foundation. R.S.E. acknowledges generous financial support from the Peter and Patricia Gruber Foundation. M.J.H. and J.C.T. thank the staff in The Doors public house for service that germinated the ideas that led to this project.
@@
All the HST data used in this paper can be found in MAST: http://dx.doi.org/10.17909/7s5v-gz68 (catalog 10.17909/7s5v-gz68).
@@ Appendix A Variability in F105W and F160W
Figures 8 and 9 report the photometric variability as a function of the mean mag in the F105W and F160W filters, as detailed in Fig. 1.